EVA Report – May 1st

EVA report #11

Date: 01st May from 4:30 pm to 6:45.

The purpose of EVA was to explore Candor Chasma and gather soil samples.

Destination: Latitude: 4251000, longitude: 52000 (UTM NAD27 CONUS)

EVA Participants: Nathan Pechon (commander), Benjamin Flasse (Medic), Chloé Peduzzi, Eléonore Lieffrig

Roads and routes per MDRS Map: Cow Dung Road 0110 (one mile towards the North), then Galileo Road 1104 (0,8 mile towards East), then Cactus Road 1104 (1 mile towards the South). Eventually we walked towards to East to Candor Chasma.

Mode of travel: Driving and hiking with Curiosity and Spirit

Narrative : After about 10 minutes driving, we arrived at Candor Chasma. We walked in the canyon for about 25 minutes. It was really beautiful, but there were a lot of trees, which is not really martian.

We saw beautiful rocks of surprising colors, such as pink, deep red or even green. Then we went back to the rovers and drove home.

EVA Report – May 1st

EVA report #10

Date: 30 April 2019 from 1pm to 3pm.

Purpose of EVA: We wanted to explore Kissing Camel since it offers an amazing point of view of all the surroundings of the station. We wanted to gather some soil samples regarding the great diversity of the local geology.

Destination: Latitude: 4249500, longitude: 518500 (UTM NAD27 CONUS)

EVA Participants: Carl-Henrik Dahlqvist(Commander), Simon Collignon(Medic), Maxime Bernard, Julien Amalaberque

Roads and routes per MDRS Map: Cow Dung Road 0110 (towards south) and we will stop between Kissing Camel Ridge West and Kissing Camel Ridge West.

Mode of travel: Driving and hiking with Spirit, Curiosity and Opportunity.

Narrative : The first challenge of this EVA was to take the rover Opportunity to Kissing Camel. It went very well since Oppy didn’t show any sign of weakness in the whole round trip. However, it was only use in "Max Range" mode and was driven very gently. I would not recommend any kind of excess regarding this rover.

Once arrived on the site, we decided to climb carefully the Kissing Camel’s hills and reach the top of one of them. We’ve received an instant reward regarding our efforts with the lovely landscapes Kissing Camel had to offer.

This point of view could also have a scientific purpose. It could be a perfect place to set up an antenna and thus reach a very large area without interfering with mountainous terrains.

After a few drone footages and a deeper exploration of the place, we took the rovers and headed back to the Hab.

Science Report – May 1st

Science Report 01 May 2019
Crew 211 – UCL to Mars
Submitted by Science Officer Maxime Bernard

Commander Carl Dahlqvist

After long late hours yesterday, our commander finally managed to make his core system work and everything is now nominal.
At the moment, he is starting his test phase with the main module and two secondary modules.

XO Simon Collignon

The RTLS (real time location system) is now working fine and he will be testing it tomorrow during an EVA.

HSO Benjamin Flasse

This morning Benjamin gathered the data from the sleep monitoring he did on Julien.
Right after that, he measured everybody’s weight, tension and brain awakeness.
In addition to that, he took measures like fat density,bone quality and water repartition intra/extra cellular using Biody Xpert like every morning.

Crew Astronomer Eleonore Lieffrig

Our Astronomer just located the quasar 3C237 on the pictures she took with the remote telescope.
She also finished the theoretical part of her work and made headway in her colloidal tracking algorithm.

Crew Engineer Julien Amalaberque

Our Engineer integrated the accelerometer and the gyroscope in his location computation in order to make it more accurate.

Crew Biologist Chloé Peduzzi

She placed her Petri dishes in the incubator and is monitoring it constantly to make sure no contamination occurred.
So far, everything seems nominal.
In the meantime, she is measuring the pH level of her Spirulina cultures.
It is increasing which is what she was expecting.

Crew Journalist Nathan Pechon

We finished to talk about the fundamental rights, freedom of speech and anti-discrimination.
The debate are still very lively and the opinions are diversified.

Crew Scientist Maxime Bernard

Keeps taking data while changing experiment conditions to find where his problem could come from.

Operations Report – May 1st

Crew 211 Operations Report 01-MAY-2019

SOL: 10

Name of person filing report: Julien Amalaberque

Non-nominal systems: The pump from the loft

Notes on non-nominal systems: The noise from the pump is beginning to slowly fade away instead of its original instant cutoff when we stop using water.

I am unsure whether it will hold for the rest of our mission or not. According to the Director a new one has been ordered.

Generator (hours run last night): 12:00

On (last night): 7:30 pm

Off (this morning): 7:30 am

Solar— SOC 72 % at 20:00 pm

Diesel Reading – 24 %

Propane Reading – 49 %

Ethanol Free Gasoline – 4 gallons

Water (auxillary tank) – 0 gallons

Water (static tank) – ~220 gallons

Auxillary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – 70 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 15 gallons

Toilet tank emptied: yes

Deimos rover used: no (out of order)

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, no

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: yes

Hours: 91.6h

Beginning charge: 100

Ending charge: 60

Currently charging: yes

Opportunity rover used: yes

Hours: 58h

Beginning charge: 100

Ending charge: 97

Currently charging: no

Curiosity rover used: yes

Hours: 93h

Beginning charge: 100

Ending charge: 70

Currently charging: yes

Notes on rovers: Nothing to report

Reason for use:

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

HabCar used and why, where? No

CrewCar used and why, where? No

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support:

Greenhab Report – May 1st

Crew 211 GreenHab Report – 1-MAY-2019

GreenHab Officer: Chloé Peduzzi

Environmental control: Ambient with door open + fan turned up to 3

Shade cloth: 80%

Average temperatures:

Low: 15°C

High: 35°C

Hours of supplemental light: 0

Daily water usage for crops: 4 gallons

Water in Blue Tank: 70 gallons

Time(s) of watering for crops: 10:30

Changes to crops: All plants look healthy. The tomato plant of the GreenHab and some basils seem a little stressed. The are more flowers on the tomato plant.

Narrative: It was also cooler today. We have harvested Basil and Radish.

Harvest: 13.9g Radish + 1.7g Basil

Support/supplies needed: I won’t have enough water for 1 week long, as it was really hot this past week, and I had to water plants accordingly.

Sol Summary – May 1st

Crew 211 Sol 10 Summary Report 1-May-2019

Sol: 10

Summary Title: At the crossroads

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: After a difficult night with the storm and a temperature drop of more than 10 °c, we made some nice pancakes to comfort us. We then all worked on our experiments as the morning EVA was postponed to the early afternoon to wait for the mud to dry. Half of the crew made an EVA to Kissing Camel from 1:00 pm to 3:00 pm. The rest of the crew left at 4:30 pm for another EVA to Condor Chasma. The objective of the first EVA was to gather information about potential places to set-up a temporary antennas system for geolocation and information transmission. This will be one of the projects for the next UCL to Mars crew for the season 2019-2020. We end the day by deploying the antenna of the vice-commander experiment and the cubelanders of the commander experiments.

Look Ahead Plan: We plan to do another EVA as the end of the stays is approaching. We will get back the antennas and the cubelanders, in the morning, before walking north following the path 1103 to the sinkhole near Hab ridge road. A second EVA is planned in the afternoon to Grey Moon. We will also try to finalize our respective experiments in order to be able to write the final report for Friday night.

Anomalies in work: No issues.

Weather: Mostly sunny, almost no wind

Crew Physical Status: Good

EVA: We had two EVAs today to enjoy the nice weather after two days of showers. The EVA to Kissing Camel allowed us to prospect for good places to setup antennas to possibly create communication and geolocation systems in order to improve the communication between the Hab and EVA teams.

Reports to be filed: Journalist Report, GreenHab Report, Operations Report, Science Report, EVA Request, EVA Report, Daily Photo Report

Support Requested: None

Crew 211 Journalist Report 01May2019

[title Journalist Report – May 1st]

Today was our tenth simulation day. The night was very cold. We woke up at 7:30 am and we had breakfast at 8:00. We try to break the monotony of our routine but the cycle is still there. A part of the crew had to go in EVA at 9 am but the ground was muddy. So the was reported on afternoon.

Before lunch, the crew worked on their experiments. For more informations read scientific report.

At 12 am, Benjamin and Eléonore prepared the lunch. We ate rest of mash potatoes fries. After meal, we wanted to discuss about Nathan’s experience and his martian constitution but Simon, Julien, Carl and Maxime went to EVA at 1 pm.

The first challenge of this EVA was to take the rover Opportunity to Kissing Camel. It went very well since Oppy didn’t show any sign of weakness in the whole round trip. However, it was only use in "Max Range" mode and was driven very gently. Simon would not recommend any kind of excess regarding this rover.Once arrived on the site, they decided to climb carefully the Kissing Camel’s hills and reach the top of one of them. They’ve received an instant reward regarding our efforts with the lovely landscapes Kissing Camel had to offer.

At 4pm, Nathan, Benjamin, Eléonore and Chloé went to EVA to explored Candor Chasma and gathered soil samples. They came back to the station at 6:30pm

After that, we prepared the Cap Com. Each of us is working on his report and is going to send it between 7 to 9 pm. We will eat rice with tomatoes.

Like usual, we will do a crew meeting to speak about the day.